# F-Bench: Rethinking Human Preference Evaluation Metrics for Benchmarking Face Generation, Customization, and Restoration

### v1

@@ 4.1 Perspective Analysis
Fig. 3 (a)-(c) demonstrates the MOS distribution across eight dimensions in the three subsets, highlighting distinct variations between dimensions. (1) In the face generation task, the Correspondence score is the highest among the three dimensions, indicating that existing generative models perform well in aligning with prompts. The distribution of Authenticity is skewed towards lower scores, suggesting that current face generation models still fall short in generating photorealistic face images that satisfy human preferences. (2) In face customization, models perform the worst in the Quality dimension overall, demonstrating the difficulty of injecting identity information without compromising image quality. Identity Fidelity shows a bimodal distribution, reflecting the dichotomous nature of human judgment regarding whether the generated face accurately represents the same identity. (3) In face restoration, models perform better in Quality than in Identity Fidelity.
@@
| Dimension | Quality | Authenticity | Correspondence | Quality | Fidelity | Correspondence | ||||||||||||
@@
| CNNIQA [30] | 0.4219 | 0.2895 | 0.3510 | 0.3037 | 0.2054 | 0.2973 | 0.2824 | 0.1931 | 0.2595 | 0.6244 | 0.4376 | 0.6620 | 0.6511 | 0.4441 | 0.6381 | 0.3765 | 0.2545 | 0.3264 |
| VGG16 [76] | 0.5846 | 0.4141 | 0.6006 | 0.5010 | 0.3508 | 0.5181 | 0.4349 | 0.2994 | 0.4723 | 0.7217 | 0.5324 | 0.7857 | 0.6735 | 0.4717 | 0.6636 | 0.5635 | 0.3932 | 0.5589 |
| VGG19 [76] | 0.5728 | 0.4043 | 0.5644 | 0.4681 | 0.3248 | 0.4479 | 0.4171 | 0.2873 | 0.4754 | 0.7952 | 0.6031 | 0.8096 | 0.7544 | 0.5417 | 0.7555 | 0.6585 | 0.4742 | 0.6465 |
| ResNet18 [21] | 0.6150 | 0.4417 | 0.6660 | 0.4600 | 0.3160 | 0.4604 | 0.4829 | 0.3360 | 0.4628 | 0.7966 | 0.6013 | 0.8006 | 0.7473 | 0.5335 | 0.7656 | 0.6308 | 0.4488 | 0.6191 |
| ResNet34 [21] | 0.6092 | 0.4354 | 0.6537 | 0.5564 | 0.3949 | 0.5401 | 0.4410 | 0.3065 | 0.5207 | 0.8141 | 0.6211 | 0.8151 | 0.7969 | 0.5878 | 0.8157 | 0.5708 | 0.4000 | 0.5810 |
| HyperIQA [78] | 0.6333 | 0.4545 | 0.6591 | 0.6106 | 0.4406 | 0.6042 | 0.4235 | 0.2912 | 0.4240 | 0.8524 | 0.6631 | 0.8205 | 0.8419 | 0.6463 | 0.8588 | 0.6628 | 0.4757 | 0.6454 |
| TReS [16] | 0.7766 | 0.5867 | 0.8047 | 0.6458 | 0.4703 | 0.6579 | 0.5708 | 0.4038 | 0.6270 | 0.8883 | 0.7083 | 0.8951 | 0.8330 | 0.6409 | 0.8658 | 0.7393 | 0.5497 | 0.7400 |
| MANIQA [101] | 0.7871 | 0.5951 | 0.8150 | 0.7308 | 0.5436 | 0.7794 | 0.6278 | 0.4443 | 0.7066 | 0.8970 | 0.7185 | 0.8952 | 0.8590 | 0.6654 | 0.8796 | 0.7849 | 0.5941 | 0.7822 |
@@
| MINTIQA [85] | 0.8312 | 0.6474 | 0.8974 | 0.8177 | 0.6306 | 0.8511 | 0.7908 | 0.5991 | 0.8667 | 0.8904 | 0.7092 | 0.8816 | 0.8524 | 0.6697 | 0.8755 | 0.8391 | 0.6413 | 0.8318 |
4.2 Model-wise Comparison
Face Generation.
Fig. 16 (a) presents the average MOS scores across three dimensions: Quality, Authenticity, and Correspondence. Correspondence scores show a steady increase over time, indicating consistent improvement in text comprehension. Quality scores generally rise with the release year, although some methods, such as Hunyuan [45] and Dreamlike [13], exhibit deviations. Hunyuan [45] oversmoothes the facial details and performs worse compared to its contemporaries, while Dreamlike [13] performs the opposite. Authenticity scores exhibit the greatest variation among models, suggesting that current face generation models place insufficient emphasis on authenticity. Except for Realistic Vision [65], SD3 [72], Flux-dev [14], and Dreamlike [13], most models lag in authenticity relative to quality, highlighting a common phenomenon of high-quality but less realistic outputs in generative models. Fig. 3 (d, e, i, j) give four examples of MOS distributions of different methods, revealing a reasonable fine-grained performance difference. More examples and analyses can be seen in supplementary materials.
Face Customization.
Fig. 16 (b) is the performance radar chart for face customization models. IP-Adapter-FaceID-Plus[28] achieves the best performance across all dimensions, followed by PhotoMaker [44]. This highlights their great robustness in identity encoders. ELITE [95] scores the lowest in all dimensions. IP-Adapter-FaceID [29] produces high-quality images but lacks accuracy in Identity Fidelity. This may be because the base model [59] introduces many AI artifacts, which affect both Identity Fidelity and Quality. InstantID [88] and FastComposer [97] score well in Identity Fidelity but underperform significantly in Quality and Correspondence, indicating strong identity preservation but limited text-driven editing capabilities. Fig. 3 (f, g, k) illustrate the model-wise MOS distributions, revealing similar trends to those observed in the radar chart.
Face Restoration.
Fig. 16 (c) displays the average performance comparison of nine face restoration models. DR2 [94] exhibits the weakest performance in both Quality and Identity fidelity. This may be due to the degradation removal process of DR2 [94], which results in the loss of fine details and excessive smoothing. In contrast, StableSR [86] and CodeFormer [107] achieve the best overall performance among the nine face restoration models, successfully preserving both intricate details and accurate identity representations. When comparing the performance on real-world and synthetic data, most methods struggle to achieve satisfactory results on complex real-world degradations, except for SAPRNet [4] and DiffBIR [49]. StableSR [86] and GFPGAN [91] show strong robustness, performing consistently well on both real-world and synthetic data.
4.3 Class-wise Comparison
For the face generation task, we analyze the Mean Opinion Score (MOS) distributions of Quality, Authenticity, and Correspondence across nine prompt categories, as shown in Fig. 5. The violin plots highlight distinct performance patterns across these categories. The Action, Expression, and Facial Attributes categories exhibit lower MOS values in all three dimensions, with notable variability. In the Action category, limited visibility of the face due to body movements affects the generation of fine facial details, resulting in reduced scores, especially in Quality and Correspondence. The Expression category poses challenges in maintaining spatial relationships and physical constraints, leading to distortion and artifacts in expressions with exaggerated facial movements, such as laughing or crying, thereby impacting both Quality and Authenticity. For Facial Attributes, generating images with precise attributes (e.g., specific eye color, hair style) is challenging, particularly when multiple attributes are specified, which lowers the scores. In contrast, the General and Background categories achieve the highest MOS scores across all dimensions, as simpler or less specific prompts are easier for the model to interpret accurately. The Clothes and Accessories categories also perform well, indicating that the model handles clothes and accessory prompts effectively. In summary, prompts with specific or complex attribute requirements tend to result in lower and more variable MOS values, while more general prompts or those with simple descriptions achieve consistently high performance across all three dimensions.
5 Performance Analysis on Current Quality Assessment Methods
5.1 Dataset and Benchmark Models
We conduct comprehensive experiments on the proposed FaceQ database to evaluate the effectiveness of current quality assessment methods, as shown in Tab.2, Tab.3, and Tab.4. The evaluation metrics for the benchmark models are SRCC, KRCC, and PLCC, which assess the correlation between the objective scores predicted by the QA methods and the subjective MOS scores provided by humans. The benchmark models are categorized into four categories:
Traditional IQA Models.
Classical Deep learning-based IQA Models.
AIGC IQA Models.
Face IQA Models.
@@ 5.2 Performance Analysis
| Dimension | Quality | Authenticity | ||||
@@
| CNNIQA [30] | 0.4661 | 0.3197 | 0.3836 | 0.4678 | 0.3216 | 0.4288 |
| ResNet18 [21] | 0.7254 | 0.5295 | 0.6982 | 0.6662 | 0.4802 | 0.6527 |
| ResNet34 [21] | 0.6693 | 0.4916 | 0.6550 | 0.5905 | 0.4170 | 0.6084 |
| VGG16 [76] | 0.7487 | 0.5513 | 0.7101 | 0.6177 | 0.4413 | 0.6381 |
| VGG19 [76] | 0.6994 | 0.5126 | 0.6722 | 0.6038 | 0.4278 | 0.5999 |
| HyperIQA [78] | 0.8269 | 0.6323 | 0.8228 | 0.7615 | 0.5585 | 0.7471 |
| MANIQA [101] | 0.8287 | 0.6403 | 0.8778 | 0.7704 | 0.5660 | 0.7619 |
| TReS [16] | 0.8656 | 0.6830 | 0.8622 | 0.7951 | 0.5879 | 0.7829 |
@@
| CNNIQA [30] | 0.3650 | 0.2506 | 0.4705 | 0.3252 | 0.2197 | 0.3112 |
| VGG16 [76] | 0.4996 | 0.3439 | 0.4269 | 0.4117 | 0.2859 | 0.4622 |
| VGG19 [76] | 0.5871 | 0.4181 | 0.5133 | 0.3054 | 0.2044 | 0.3594 |
| ResNet18 [21] | 0.5977 | 0.4221 | 0.5722 | 0.5554 | 0.3909 | 0.6161 |
| ResNet34 [21] | 0.6259 | 0.4475 | 0.6168 | 0.5036 | 0.3539 | 0.6136 |
| HyperIQA [78] | 0.7926 | 0.5965 | 0.7926 | 0.7783 | 0.5845 | 0.8163 |
| MANIQA [101] | 0.8356 | 0.6482 | 0.8467 | 0.7597 | 0.5655 | 0.8081 |
| TReS [16] | 0.8287 | 0.6391 | 0.8501 | 0.7287 | 0.5441 | 0.8009 |
@@
Traditional IQA methods, including both NR-IQA and FR-IQA models, generally perform poorly across all subsets and dimensions, with low SRCC, KRCC, and PLCC scores. These methods exhibit a significant discrepancy from human perception, particularly in quality and authenticity evaluation, likely due to their limited ability to handle generative artifacts such as background blur.
@@
Deep learning-based IQA models, such as MANIQA [101] and HyperIQA [78], outperform traditional models with considerably higher correlation scores across dimensions. These models demonstrate enhanced sensitivity to image quality attributes relevant to human perception, achieving higher alignment with subjective ratings. However, they still encounter challenges with specific dimensions, especially when assessing complex attributes.
@@
AIGC-focused models like MINTIQA [85] show competitive performance, especially in the quality and correspondence dimensions on generative tasks. MINTIQA [85] performs consistently well on synthetic and real-world subsets, suggesting it is better adapted to handle generative content. In contrast, CLIPScore [24] and BLIPScore [40] show weaker correlations, likely due to their focus on high-level semantics rather than low-level details crucial for image quality assessment.
@@
Face-specific IQA models (e.g., ArcFace [12], SER-FIQ [80], DSL-FIQA [8]) excel in evaluating ID fidelity, as seen in the real-world subset, where DSL-FIQA achieves high SRCC and PLCC scores. These models are optimized for identity consistency, making them particularly effective for tasks involving face quality and fidelity, though they may lack robustness in other general quality dimensions.

### v2

@@ 4 F-Bench: Benchmarking Current Generative Face Models
In this section, we establish a human preference benchmark F-Bench for today’s state-of-the-art face generation, customization, and restoration methods, from quality, authenticity, ID fidelity, and correspondence dimensions. This benchmark analyzes the strengths and weaknesses, providing new insights and potential directions for future work.
@@ 4.1 Perspective Analysis
Fig.3 demonstrate the MOS distribution across eight dimensions in the three subsets, highlighting distinct variations between dimensions. (1) In the face generation task, the Correspondence score is the highest among the three dimensions, indicating that existing generative models perform well in aligning with prompts. The distribution of Authenticity is skewed towards lower scores, suggesting that current face generation models still fall short in generating photorealistic face images that satisfy human preferences. (2) In face customization, models perform the worst in the Quality dimension overall, demonstrating the difficulty of injecting identity information without compromising image quality. Identity Fidelity shows a bimodal distribution, reflecting the dichotomous nature of human judgment regarding whether the generated face accurately represents the same identity. (3) In face restoration, models perform better in Quality than in Identity Fidelity.
4.2 Model-wise Comparison
Face Generation. Fig.17a presents the average MOS scores across three dimensions: Quality, Authenticity, and Correspondence. Correspondence scores show a steady increase over time, indicating consistent improvement in text comprehension. Quality scores generally rise with the release year, although Hunyuan [58] and Dreamlike [2] exhibit deviations. Hunyuan [58] oversmoothes the facial details and performs worse compared to its contemporaries, while Dreamlike [2] performs the opposite. Authenticity scores exhibit the greatest variation among models, suggesting that current face generation models place insufficient emphasis on authenticity. Except for Realistic Vision [8], SD3 [9], Flux-dev [3], and Dreamlike [2], most models lag in authenticity relative to quality, highlighting a common phenomenon of high-quality but less realistic outputs in generative models. Fig.3d,e give some examples of MOS distributions of different methods, revealing a reasonable fine-grained performance difference.
Face Customization. Fig.17b is the performance radar chart for face customization models. IP-Adapter-FaceID-Plus[5] achieves the best performance across all dimensions, followed by PhotoMaker [57]. This highlights their great robustness in identity encoders. ELITE [101] scores the lowest in all dimensions. IP-Adapter-FaceID [4] produces high-quality images but lacks accuracy in Identity Fidelity. This may be because the base model [71] introduces many AI artifacts, which affect both Identity Fidelity and Quality. InstantID [94] and FastComposer [104] score well in Identity Fidelity but underperform significantly in Quality and Correspondence, indicating strong identity preservation but limited text-driven editing capabilities. Fig.3h illustrate the model-wise MOS distributions, revealing similar trends to those observed in the radar chart.
Face Restoration. Fig.17c displays the average performance comparison of nine face restoration models. DR2 [100] exhibits the weakest performance in both Quality and Identity fidelity. This may be due to the degradation removal process of DR2 [100], which results in the loss of fine details and excessive smoothing. In contrast, StableSR [92] and CodeFormer [116] achieve the best overall performance among the nine face restoration models, successfully preserving both intricate details and accurate identity representations. When comparing the performance on real-world and synthetic data, most methods struggle to achieve satisfactory results on complex real-world degradations, except for SAPRNet [15] and DiffBIR [62]. StableSR [92] and GFPGAN [97] show strong robustness, performing consistently well on both real-world and synthetic data.
4.3 Class-wise Comparison
For the face generation task, we analyze the Mean Opinion Score (MOS) distributions of Quality, Authenticity, and Correspondence across nine prompt categories, as shown in Fig.5. The violin plots highlight distinct performance patterns across these categories. The Action, Expression, and Facial Attributes categories exhibit lower MOS values in all three dimensions, with notable variability. In the Action category, limited visibility of the face due to body movements affects the generation of fine facial details, resulting in reduced scores, especially in Quality and Correspondence. The Expression category poses challenges in maintaining spatial relationships and physical constraints, leading to distortion and artifacts in expressions with exaggerated facial movements, such as laughing or crying, thereby impacting both Quality and Authenticity. For Facial Attributes, generating images with precise attributes (e.g., specific eye color, hair style) is challenging, particularly when multiple attributes are specified, which lowers the scores. In contrast, the General and Background categories achieve the highest MOS scores across all dimensions, as simpler or less specific prompts are easier for the model to interpret accurately. The Clothes and Accessories categories also perform well, indicating that the model handles clothes and accessory prompts effectively. Overall, prompts with specific or complex attribute requirements tend to result in lower MOS scores, while general or simple prompts yield consistently high scores across dimensions. This can be attributed to the absence of high-quality fine-grained text-face pair datasets during the training of generative models.
@@ 4.3 Class-wise Comparison
| Dimension | Quality | Authenticity | Correspondence | Quality | ID Fidelity | Correspondence | ||||||||||||
@@
| IFQA[41] | 0.2313 | 0.1549 | 0.3261 | 0.1597 | 0.1062 | 0.2097 | 0.1597 | 0.1077 | 0.2260 | 0.2401 | 0.1614 | 0.2220 | 0.1503 | 0.0973 | 0.1739 | 0.3731 | 0.2546 | 0.3941 |
| CNNIQA* [42] | 0.4219 | 0.2895 | 0.3510 | 0.3037 | 0.2054 | 0.2973 | 0.2824 | 0.1931 | 0.2595 | 0.6244 | 0.4376 | 0.6620 | 0.6511 | 0.4441 | 0.6381 | 0.3765 | 0.2545 | 0.3264 |
| VGG16* [81] | 0.5846 | 0.4141 | 0.6006 | 0.5010 | 0.3508 | 0.5181 | 0.4349 | 0.2994 | 0.4723 | 0.7217 | 0.5324 | 0.7857 | 0.6735 | 0.4717 | 0.6636 | 0.5635 | 0.3932 | 0.5589 |
| VGG19* [81] | 0.5728 | 0.4043 | 0.5644 | 0.4681 | 0.3248 | 0.4479 | 0.4171 | 0.2873 | 0.4754 | 0.7952 | 0.6031 | 0.8096 | 0.7544 | 0.5417 | 0.7555 | 0.6585 | 0.4742 | 0.6465 |
| ResNet18* [29] | 0.6150 | 0.4417 | 0.6660 | 0.4600 | 0.3160 | 0.4604 | 0.4829 | 0.3360 | 0.4628 | 0.7966 | 0.6013 | 0.8006 | 0.7473 | 0.5335 | 0.7656 | 0.6308 | 0.4488 | 0.6191 |
| ResNet34* [29] | 0.6092 | 0.4354 | 0.6537 | 0.5564 | 0.3949 | 0.5401 | 0.4410 | 0.3065 | 0.5207 | 0.8141 | 0.6211 | 0.8151 | 0.7969 | 0.5878 | 0.8157 | 0.5708 | 0.4000 | 0.5810 |
| HyperIQA* [83] | 0.6333 | 0.4545 | 0.6591 | 0.6106 | 0.4406 | 0.6042 | 0.4235 | 0.2912 | 0.4240 | 0.8524 | 0.6631 | 0.8205 | 0.8419 | 0.6463 | 0.8588 | 0.6628 | 0.4757 | 0.6454 |
| TReS* [25] | 0.7766 | 0.5867 | 0.8047 | 0.6458 | 0.4703 | 0.6579 | 0.5708 | 0.4038 | 0.6270 | 0.8883 | 0.7083 | 0.8951 | 0.8330 | 0.6409 | 0.8658 | 0.7393 | 0.5497 | 0.7400 |
| MANIQA* [109] | 0.7871 | 0.5951 | 0.8150 | 0.7308 | 0.5436 | 0.7794 | 0.6278 | 0.4443 | 0.7066 | 0.8970 | 0.7185 | 0.8952 | 0.8590 | 0.6654 | 0.8796 | 0.7849 | 0.5941 | 0.7822 |
| PromptIQA* [14] | 0.7823 | 0.5907 | 0.8245 | 0.7232 | 0.5327 | 0.7558 | 0.6514 | 0.4698 | 0.7065 | 0.9071 | 0.7315 | 0.8821 | 0.8693 | 0.6792 | 0.8876 | 0.8348 | 0.6342 | 0.8281 |
@@
| MINTIQA[91] | 0.7002 | 0.5160 | 0.8047 | 0.6217 | 0.4450 | 0.6826 | 0.5491 | 0.3891 | 0.6584 | 0.5378 | 0.3775 | 0.5574 | 0.4088 | 0.2816 | 0.4041 | 0.2886 | 0.1932 | 0.2959 |
| MINTIQA* [91] | 0.8312 | 0.6474 | 0.8974 | 0.8177 | 0.6306 | 0.8511 | 0.7908 | 0.5991 | 0.8667 | 0.8904 | 0.7092 | 0.8816 | 0.8524 | 0.6697 | 0.8755 | 0.8391 | 0.6413 | 0.8318 |
| F-Eval (Ours) | 0.8486 | 0.6670 | 0.9085 | 0.8312 | 0.6585 | 0.8578 | 0.8471 | 0.6637 | 0.9106 | 0.9462 | 0.7961 | 0.9461 | 0.9188 | 0.7640 | 0.9322 | 0.9460 | 0.7959 | 0.9457 |
5 F-Eval Method
In this section, we introduce our one-for-all generated face quality assessment method, F-Eval, towards predicting quality scores from multiple dimensions for three tasks using one model.
5.1 Overall Architecture
The overall architecture of F-Eval is depicted in Fig. 6. As shown on the left, F-Eval can assess four dimensions: quality, authenticity, text-image correspondence, and identity fidelity, with multimodal inputs, i.e.text prompts, image and optional reference image, to output quality-related answers. On the right, the input image(s) can originate from all three tasks. F-Eval employs a vision encoder and a face encoder to extract visual features, with projectors projecting both into a unified language space. User prompts are encoded into textual embeddings through a tokenizer. These multimodal features are concatenated and fed into a pre-trained large language model for the final output.
5.2 Model Design
Vision Encoding.
For input image(s), F-Eval adopt a vision encoder to extract global features and a face encoder captures face features related to facial structure and identity. Specifically, the vision encoder, Qwen-VL-2.5-ViT [11], is tuned on FaceQ to adjust different dimensions. Face encoder is Arcface [22], which is frozen. Two two-layer multi-layer perception projectors then map the features .
Feature Fusion via the LLM.
For input prompts, it is tokenized into text tokens and then concatenated with visual tokens. Here, Qwen-VL2.5-7B [11] is employed as the pre-trained LLM backbone to combine visual tokens and text tokens to perform multi-modal learning.. Finally, the output features of the LLM are decoded with a text decoder and then projected to the prompt-corresponded quality space.
5.3 Multi-Expert LoRA Tuning
Instruction Tuning.
Previous IQA methods are often designed for specific dimensions [109, 115], addressing only one aspect of the problem or requiring retraining of model weights for different dimensions. Leveraging the generalization ability of LLMs, instruction tuning allows for the integration of different dimensions within a unified framework. Therefore, we construct diverse instruction data for the four facial quality evaluation dimensions in the Face-Q dataset and fine-tune F-Eval by adjusting the visual encoder, projector, and LLM parameters to adapt it to the specific task. As a result, F-Eval is capable of evaluating facial quality using unified weights.
Mixture of LoRA Experts (MoLE).
Tuning a holistic quality model for different dimensions using a single LoRA expert introduces inherent interference between dimensions. Hence, F-Eval integrates four LoRA experts for four dimensions when tuning vision encoder and LLM, constituting MoLE modules. A router is trained to classify input prompts into the four dimensions and output the corresponding dimension ID . Specifically, for a frozen linear layer with input and weight matrix , a LoRA expert is trained to learn a low-rank decomposed update:
where and are the low-rank matrices, is the rank, and is scaling factor. For prompt of certain dimension, the corresponding expert is activated based on the dimension ID , which can be formulated as:
where is the k-th LoRA expert.
6 Experiments
6.1 Implementation Details
The LLM used in F-Eval is Qwen-2.5VL [11], with an input token channel dimension of 3584. The vision encoder, face encoder, and LLM are all frozen during training. The ranks of LoRA experts are 16,32,16,16, respectively. The LoRA alphas are set equal to LoRA ranks. LoRA dropout is set as 0.05. The model is trained on 2 NVIDIA GeForce RTX 3090 GPUs with flash-attn [21] enabled. We use AdamW Optimizer to train the network for 10 epochs with a batch size of 4 and learning rate .
6.2 Comparison with existing QA Models
We conduct comprehensive experiments on the proposed FaceQ database to evaluate the effectiveness of both current quality assessment methods and our proposed methods, as shown in Tab.2, Tab.3, and Tab.4. The evaluation metrics are Spearman rank-order correlation coefficient (SRCC), Kendall rank order correlation coefficient (KRCC), and Pearson linear correlation coefficient (PLCC), which assess the correlation between predicted scores by the QA methods and the MOS scores provided by humans. Higher metrics indicated better alignment with human preferences. The comparison QA models falls into four categories: 1) Traditional IQA Models. NR-IQA models include NIQE [67], ILNIQE [114], HOSA [105], BPRI-PSS [64], FISBLIM [27]. FR-IQA models include LPIPS [115], SSIM [99] and PSNR [34]. 2) Classical Deep learning-based IQA Models. This category includes CNNIQA [42], VGG16 [81], VGG19 [81], ResNet18 [29], ResNet34 [29], HyperIQA [83], TReS [25], MANIQA [109],PromptIQA[14]. All models are retrained on FaceQ with an 80:20 training/testing split. 3) AIGC IQA Models. This category includes pre-trained vision-language models, including CLIPScore [32], BLIPScore [53], zero-shot metric ImageReward [106], and the AIGC-specific QA model MINTIQA [91]. 4) Face IQA Models. This category includes SER-FIQ [85], DSL-FIQA [18], ArcFace [22], IFQA[41].
@@ 6.3 Performance Analysis
| Dimension | Quality | ID Fidelity | ||||
@@
| IFQA [41] | 0.2273 | 0.1527 | 0.3348 | 0.3214 | 0.2185 | 0.3874 |
| CNNIQA* [42] | 0.4661 | 0.3197 | 0.3836 | 0.4678 | 0.3216 | 0.4288 |
| ResNet18* [29] | 0.7254 | 0.5295 | 0.6982 | 0.6662 | 0.4802 | 0.6527 |
| ResNet34*[29] | 0.6693 | 0.4916 | 0.6550 | 0.5905 | 0.4170 | 0.6084 |
| VGG16* [81] | 0.7487 | 0.5513 | 0.7101 | 0.6177 | 0.4413 | 0.6381 |
| VGG19* [81] | 0.6994 | 0.5126 | 0.6722 | 0.6038 | 0.4278 | 0.5999 |
| HyperIQA* [83] | 0.8269 | 0.6323 | 0.8228 | 0.7615 | 0.5585 | 0.7471 |
| MANIQA* [109] | 0.8287 | 0.6403 | 0.8778 | 0.7704 | 0.5660 | 0.7619 |
| TReS* [25] | 0.8656 | 0.6830 | 0.8622 | 0.7951 | 0.5879 | 0.7829 |
| PromptIQA* [14] | 0.8519 | 0.6661 | 0.8733 | 0.8220 | 0.6275 | 0.8254 |
| F-Eval (Ours) | 0.8692 | 0.6855 | 0.9009 | 0.8507 | 0.6731 | 0.8726 |
@@
| IFQA [41] | 0.0251 | 0.0210 | 0.1221 | 0.0893 | 0.0603 | 0.2086 |
| CNNIQA* [42] | 0.3650 | 0.2506 | 0.4705 | 0.3252 | 0.2197 | 0.3112 |
| VGG16*[81] | 0.4996 | 0.3439 | 0.4269 | 0.4117 | 0.2859 | 0.4622 |
| VGG19* [81] | 0.5871 | 0.4181 | 0.5133 | 0.3054 | 0.2044 | 0.3594 |
| ResNet18* [29] | 0.5977 | 0.4221 | 0.5722 | 0.5554 | 0.3909 | 0.6161 |
| ResNet34* [29] | 0.6259 | 0.4475 | 0.6168 | 0.5036 | 0.3539 | 0.6136 |
| HyperIQA* [83] | 0.7926 | 0.5965 | 0.7926 | 0.7783 | 0.5845 | 0.8163 |
| MANIQA* [109] | 0.8356 | 0.6482 | 0.8467 | 0.7597 | 0.5655 | 0.8081 |
| TReS* [25] | 0.8287 | 0.6391 | 0.8501 | 0.7287 | 0.5441 | 0.8009 |
| PromptIQA* [14] | 0.8115 | 0.6186 | 0.8645 | 0.7398 | 0.5526 | 0.8297 |
| F-Eval(Ours) | 0.8448 | 0.6577 | 0.8705 | 0.7957 | 0.6057 | 0.8366 |
@@
Traditional IQA methods, including both NR-IQA and FR-IQA models, generally perform poorly across all tasks and dimensions, with low SRCC, KRCC, and PLCC scores.
@@
CLIPScore [32] BLIPScore [53], and ImageReward [106] exhibit weak correlations with human preference, likely due to their focus on high-level semantics rather than low-level details crucial for image quality assessment. MINTIQA [91] achieves best performance among current AIGC IQA methods, but its performance is still unsatisfactory, particularly on the customization task. Fine-tuning on our dataset leads to significant improvements, highlighting the great potential of FaceQ.
@@
ArcFace similarity [22] demonstrates its ability to extract identity-related features. However, it lacks perceptual sensitivity for image quality, leading to poor performance in quality assessment. DSL-FIQA [18], as the latest method, achieves the best performance in predicting quality scores among face IQA methods, but it is still far from satisfactory. SER-FIQ fails to align with human perception across most dimensions and tasks. Experiments confirm that these natural face evaluation algorithms struggle to assess the complex characteristics of AIGFs, highlighting the need for an effective AIGF evaluator.
F-Eval (Ours).
Our method achieves optimal performance across all tasks and dimensions, with the most significant gain in the customization task, as existing methods rarely focus human preferences in face customization. Among the four dimensions, the improvements in correspondence and ID fidelity stand out, thanks to dual-stream vision encoding and the powerful LLM backbone for text understanding and modality fusion. The MoLE technique allows the model to learn distinct features for each dimension.
6.4 Ablation Study
We conducted the ablation study to validate the utility of the core components of F-Eval in Tab. 5.
Effectiveness of Vision Encoding.
1) The effectiveness of the face encoder is demonstrated by comparing the second and last rows. Without the face encoder, a significant performance drop is observed, indicating that the extracted features help the LLM interpret face images. 2) Removing the projector leads to a dramatic performance drop.
Effectiveness of MoLE.
We conducted three sets of LoRA adaptation experiments: fine-tuning a single LoRA for all dimensions and tasks, fine-tuning three LoRA experts for three tasks, and fine-tuning four LoRA experts for four dimensions. The results show that fine-tuning multiple dimensions with a single LoRA leads to performance deterioration, which further confirms the independence of the characteristics across different dimensions.
| Feature & Strategy | Face Generation | Face Customization | Face Restoration syn | Face Restoration rw | |||||||||||
| Vision | Projector | Face | Single LoRA | Task LoRA | Dim. LoRA | Quality | Authen. | Corres. | Quality | ID. | Corres. | Quality | ID. | Quality | ID. |
| ✔ | ✔ | ✔ | 0.808 | 0.750 | 0.846 | 0.924 | 0.843 | 0.927 | 0.777 | 0.776 | 0.707 | 0.478 | |||
| ✔ | ✔ | ✔ | 0.815 | 0.752 | 0.835 | 0.924 | 0.855 | 0.921 | 0.812 | 0.783 | 0.716 | 0.603 | |||
| ✔ | ✔ | ✔ | ✔ | 0.793 | 0.776 | 0.819 | 0.941 | 0.885 | 0.942 | 0.838 | 0.823 | 0.802 | 0.671 | ||
| ✔ | ✔ | ✔ | ✔ | 0.823 | 0.780 | 0.838 | 0.942 | 0.873 | 0.942 | 0.824 | 0.826 | 0.754 | 0.638 | ||
| ✔ | ✔ | ✔ | ✔ | 0.849 | 0.831 | 0.847 | 0.946 | 0.919 | 0.946 | 0.869 | 0.851 | 0.845 | 0.796 | ||